Find the surface area, to the nearest tenth, of a cylinder with diameter 5 cm and height 8 cm. (S=2(pi)rh+2(pi)r^2)
Thank you so, so much for all your help!!!!

Answer by nerdybill(7384)

Find the surface area, to the nearest tenth, of a cylinder with diameter 5 cm and height 8 cm. (S=2(pi)rh+2(pi)r^2)
.
Assuming pi is 3.14
radius = diameter/2 = 5/2 = 2.5 cm
.
simply plug in the given dimension into:
S=2(pi)rh+2(pi)r^2
S=2(pi)r(h+r)
S=2(3.14)(2.5)(8+2.5)
S=2(3.14)(2.5)(10.5)
S= 164.85
Rounding to nearest tenth:
S = 164.9 square cm